60% of students in college play sports, 14% of the students who play sports play baseball. what percent of the students in the college play baseball?
1
Expert's answer
2013-03-29T14:05:36-0400
Let X be the total number of students.
Then the number of students which play sports is equal to
0.6 * X
Therefore the number of students who play baseball is
0.14 * 0.6 * X = 0.084 * X.
Hence
100*0.084 = 8.4%
percent of the students in the college play baseball.